OpenStudy (anonymous):

i need help solving this: x^(2/3)-2x^(1/3)=15. And the intrsutions say explain and demonstrate how to use substitution to solve the equation. well i don't understand this at all. please help!

OpenStudy (anonymous):

treat x^(2/3) as x^(1/3)^2

OpenStudy (anonymous):

ok put \[z=x^{\frac{1}{3}}\]

OpenStudy (anonymous):

then factor as x^2-2x-15=0, and substitute x^1/3 back in for x

OpenStudy (anonymous):

then by the laws of exponents \[z^2=(x^{\frac{1}{3}})^2=x^{\frac{2}{3}}\]

OpenStudy (anonymous):

so by replacing \[x^{\frac{1}{3}}\] by \[z\] you get an equation that looks like \[z^2-2z=15\]

OpenStudy (anonymous):

so the answer should be 5?

OpenStudy (anonymous):

there are two answers

OpenStudy (anonymous):

and you can solve this for z by writing \[z^2-2z-15=0\] \[(z-5)(z+3)=0\] so \[z=5\] or \[z=-3\]

OpenStudy (anonymous):

that is what i was getting 5 or -3 but then when i try to substitute back in i do not get anything that looks right. i absolutely think that i am doing something wrong still

OpenStudy (anonymous):

but now we have to remember that \[z=x^{\frac{1}{3}}\]so that means that \[x^{\frac{1}{3}}=5\] or \[x^{\frac{1}{3}}=-3\]

OpenStudy (anonymous):

and so \[x=(5)^2=125\] or \[x=(-3)^3=-27\]

OpenStudy (anonymous):

so i substitute back in -25 or 127?

OpenStudy (anonymous):

i mean -27 or 125

OpenStudy (anonymous):

you substitute x^(1/3) back in for z

OpenStudy (anonymous):

x^(1/3)=5 x^(1/3)=-3

OpenStudy (anonymous):

then cube both sides of each equation to find x

OpenStudy (anonymous):

(x^(1/3))^3=5^3 x=125 (x^(1/3))^3=-3^3 x=-27

OpenStudy (anonymous):

so both are good solutions? im trying to follow sorry, i am just very bad at this

OpenStudy (anonymous):

so both are good solutions? im trying to follow sorry, i am just very bad at this

OpenStudy (anonymous):

Yes both are good solutions because they are in the domain of the original function.